OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 4x2 + 6x + 2 = 0

OpenStudy (anonymous):

@thomaster @Compassionate

OpenStudy (anonymous):

Numerical Answers Expected!

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@hoblos

OpenStudy (anonymous):

they can probably help with this question

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

b^2-4ac

OpenStudy (anonymous):

is that numerical value

OpenStudy (anonymous):

if that is a perfect square or 0 then we can factor.

OpenStudy (anonymous):

but it is not favorable

OpenStudy (anonymous):

6^2-4*2*4=36-32=4=2^2

OpenStudy (anonymous):

otherwise to factor ax^2+bx+c=0 find a*c b=p+q p*q=ac

OpenStudy (anonymous):

Your in connexus right?

OpenStudy (anonymous):

so it would be 4?